INAF queue observing, 2016 Jan 13-14

Observer: R. Carini, A. Marchetti
Telescope Operator: G. Bechetti
Support Astronomer: Barry Rothberg
Instrument: LBC, MODS/PARALLEL

Summary

The INAF time is only up to 8.30 UT (first half of the night)

At the sunset the sky is clear, the humidity is about 20%

Night duration (12\xB0 el twilight) 19:00-6:00 ( 02:00am-01:00pm UT) = 11 h 00m

Observing time: 11 hrs 0 m (100%)

Weather time loss: 0 hrs 00m

technical time loss 0 m

Data Summary

Project_RUN Objects Filter ExpTime Notes
SKY FLAT   U_spec, F972N20   just three images are good for F972N20
Saracco M004XLSms2 red grating 1h
Nascimbeni GJ3470 Uspec ,F972N20 4 h completed

Details

00:09 Take the test images for MODS * mods1r.20160114.0001, mods1b.20160114.0001

00:37 Take sky flats for LBC for Uspec and F972N20

01:10 There are problems with the secondary mirror

01:50 problem resolved, the TO sends the preset e begins the collimation

02:05 we send the acquisition of Saracco acqMODS mods.4.XLSms2.acq
  • 02:11 problem with the preset, we can not do the acquisition.
  • TO stop the IIF and GSI, restart the TCS
  • 2:16 acqMODS mods.4.XLSms2.acq mods1r.0002-0007
  • 01:41 execMODS mods.4.XLSms2_4.obs (integrated the last two frames of the ob) mods1r.0008-0009
  • seeing about 1.2 (peaks to 1.4'')
  • 02:52 seeing 1.4' on guider, ut 1.1''on dimm
  • a drop in temperature could be caused the worsening and variability of the seeing in the last hour
  • 02:57 seeing 1.3'' on guider, 1.2 on the dimm
  • 03:01 seein 1.4 on the guider and 1.2'' on the dimm
  • 03:13 seeing on the dimm 1'''

03:18 execMODS mods.4.XLSms2_5.obs mods1r.0010-0011
  • seeing about 1.1'', during the integration the seeing get better, to 0.7'''

3:59 slew to the Nascimbeni Program (Transit in parallel mode)
  • acqMODS mods.2.GJ3470.mms mods1r.0012-0014
  • 04:04 dofpia
  • problem with the collimation, the TO cleans the adaptive optics in the red side
  • 04:20 dofpia, the collimation doesn't works, there is a huge coma, maybe it is due a the presence of a gradient of temperature in the mirror
  • 04:27 TO increases Z8 and run dofpia again

04:41 run the science with MODS execMODS mods.2.GJ3470.obs mods1r.0015-000082, mods1b.0041-
  • defocus with LBCred
  • 4:45 start the integration with LBC GJ3470_f972.ob, the star have counts more than 30.000, we stop the integration to defocus
  • 4:51 start again the integration with LBC GJ3470_f972.ob
  • 4.57 defocus a little bit more
  • seeing 1.2''
  • 05:35 defocus, seeing about 1.1''
  • 06:03 seeing ~ 0.8"
  • 06.24 seeing between 0.7 and 0.8"

07:27 MIRROR PANIC. during the integration the TO cleans the optics, this doesn't damage integration with MODS, but the LBC images loose the focus, the images are a big coma.
  • we stop and run dofpia
  • 7:38 run again the ob GJ3470_f972.ob
  • there is still coma in the images 073840 and 074016, after the corrections the image are again good, but the count are between 10000 and 15000. The AS suggest to do not focus, we risk to have the coma again in the image.
  • seeing about 0.76'''

08:33 stop the integration, the transit is finished

-- %USERSIG{RobertaCarini - 2016-01-13}%

Comments

 
Topic revision: r5 - 14 Jan 2016, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback